# Rethinking Analytical Processing in the GPU Era

### v2

@@ Abstract.
The era of GPU-powered data analytics has arrived. In this paper, we argue that recent advances in hardware (e.g., larger GPU memory, faster interconnect and IO, and declining cost) and software (e.g., composable data systems and mature libraries) have removed the key barriers that have limited the wider adoption of GPU data analytics. We present Sirius, a prototype open-source GPU-native SQL engine that offers drop-in acceleration for diverse data systems. Sirius treats GPU as the primary engine and leverages libraries like libcudf for high-performance relational operators. It provides drop-in acceleration for existing databases by leveraging the standard Substrait query representation, replacing the CPU engine without changing the user-facing interface. On TPC-H, Sirius achieves 7 speedup when integrated with DuckDB in a single node at the same hardware rental cost, and up to 12.5 speedup when integrated with Apache Doris in a distributed setting. 00footnotetext: ∗Equal contributions
1. Introduction
@@
| Rental Cost | $7.344/h (AWS) | $3.2/h (Lambda Labs) |
Despite their great potential, however, GPU-based SQL engines (omn, 2025; kin, 2025; the, 2025; cud, 2025; spa, 2025) have not yet seen mainstream adoption. Prior efforts, both academic and commercial, have been constrained by the following four key challenges and thereby remain niche solutions.
@@
In this paper, we argue that the hardware and software foundations are finally in place today to overcome these challenges and enable scalable GPU data analytics. On the hardware side, GPU memory capacity doubles almost every generation, starting from Volta (32 GB), to Ampere (80 GB), Hopper (192 GB), and most recently Blackwell (288 GB). Better interconnects such as PCIe Gen6, NVLink-C2C (nvl, 2025), and GPU Direct (gpu, 2025) significantly reduce the data movement overhead between GPU and other system components. This allows a GPU to process data beyond on-device memory with very fast speed, enabling TBs of analytics even on a single GPU and more with distribution. Meanwhile, GPUs are increasingly affordable and accessible, especially for older generations which are already sufficient for data analytics workloads.
On the software side, the GPU ecosystem has significantly matured. Libraries such as libcudf (cud, 2025) provides high-performance primitives like joins and aggregations that a GPU SQL engine can build upon. The rise of composable data systems (Pedreira et al., 2023)—driven by open standards like Substrait (sub, 2025) for query representation, Apache Arrow and Parquet for columnar data formats—enable greater interoperability. A modern GPU engine can reuse existing components, such as SQL parser and query optimizer, to drastically reduce the complexity of building the system from the ground up.
To demonstrate the feasibility of GPU-native databases, we have built Sirius111https://github.com/sirius-db/sirius, a prototype open-source SQL engine that uses GPU as the primary execution device and delivers drop-in acceleration across diverse data systems. Sirius integrates seamlessly with existing database systems via the standard Substrait query representation. For example, plugging Sirius into DuckDB causes minimal change to the user-facing interface. Instead of executing the query using DuckDB’s CPU engine, the Substrait plan is routed to Sirius for GPU-native execution. Sirius builds on GPU libraries such as libcudf (cud, 2025), RMM (rmm, 2025), and NCCL (ncc, 2025), reusing optimized implementations of core relational operators like joins, filters, aggregations, and data shuffle. Thanks to its modular design, Sirius also allows developers to easily switch the operator implementation between these GPU libraries and custom CUDA kernels. This flexibility reduces engineering overhead while enabling Sirius to achieve end-to-end GPU acceleration. Our detailed evaluation on TPC-H benchmark shows that Sirius can deliver 7 speedup as a drop-in accelerator for DuckDB (Raasveldt and Mühleisen, 2019) (single-node) and up to 12.5 speedup as a drop-in accelerator for Apache Doris (dor, 2025) (distributed).
@@ 2.1. Hardware Trend
PCIe, the interconnect between GPU and CPU, has traditionally been a severe performance bottleneck due to its relatively low bandwidth. However, PCIe bandwidth has been doubling every two years with PCIe Gen6 offering 128 GB/s, which is comparable to CPU memory bandwidth. The NVLink-C2C (nvl, 2025) technology takes this even further, enabling 900 GB/s unidirectional bandwidth between CPU and GPU. In a GH200 superchip, the Hopper GPU can access host memory at more than 400 GB/s, which is more than the maximum bandwidth between the CPU and main memory. These faster interconnects allow a GPU analytics engine to support data larger than the GPU device memory, enabling high-speed access to terabytes of main memory and beyond.
Faster Network and Storage. Besides faster memory speed, modern GPUs also have increasingly better support for network and storage, allowing a GPU engine to go beyond in-memory data processing. Technologies like GPUDirect Storage and GPUDirect RDMA allow GPUs to access data directly from storage or network with minimal CPU involvement at very high speed. Similar to interconnect technology, storage and network bandwidth have also grown rapidly in recent years—and are expected to continue improving (see Figure 1(c) and 1(d)). Moreover, the recent S3 over RDMA technique further enables high-speed object store for GPUs, which can achieve storage access speed at 200 GB/s (s3r, 2025). With significantly higher computational power, GPUs are better positioned than CPUs to fully utilize these high-bandwidth resources.
Declining GPU Cost. While GPUs of the latest generation remain in high demand with high cost, older generations are becoming increasingly accessible in the cloud at substantially lower cost. For instance, the GH200 instance shown in Table 1 has an on-demand price of $3.2/hour in Lambda Labs, which is cheaper than many high-end CPU instances. The on-demand H100 prices in many GPU cloud providers has seen significant price drop from $8/hour in March 2023 to around $3/hour in 2025 (h10, 2025a, b). Even in major cloud provider such as AWS, in 2025 alone, the A100, H100, and H200 on-demand price have dropped by 33%, 44%, and 25% respectively (aws, 2025).
@@ 2.2. Software Trend
Composable Data Systems. Modern open-source data systems are embracing composability (Pedreira et al., 2023) (e.g. Velox (Pedreira et al., 2022)), where different system components can be independently developed and interoperate in a data system. A new database can reuse existing components such as the language frontend, query optimizer, or execution engine instead of reinventing the wheel. A crucial component in these systems is Substrait (sub, 2025), a unified IR for physical and logical representations of query plans. The trend of composable data systems significantly lowers the barrier to build a GPU database—only the execution engine needs to be developed from the ground up tailored for GPU but other system components could be reused.
Maturing GPU Libraries. Open-source GPU libraries optimized for data processing have emerged and are rapidly reaching maturity in recent years. For example, libcudf (cud, 2025) provides GPU-efficient implementation for relational operators such as joins, aggregations, and sorting. RMM (rmm, 2025) complements this by offering a high-performance GPU memory allocator, serving as a foundation for GPU-centric buffer managers. On the distributed side, the NCCL (ncc, 2025) provides high-speed communication between GPU devices over various interconnect, such as PCIe, NVLink, and Ethernet/RDMA.
@@ 3.1. Design Overview
Taking this into account, we design Sirius as a GPU-native engine: it treats GPUs as the primary execution engine, aiming to run the entire query plan—from scan to result—on the GPU. Sirius differs from systems that retrofit GPU acceleration onto traditional CPU-optimized engines while maintaining backward compatibility (spa, 2025; vel, 2025), or hybrid systems that split execution between CPU and GPU (Yogatama et al., 2022; Kabić et al., 2025; Yogatama et al., 2024; Chrysogelos et al., 2019). Instead, it builds its execution model, buffer management, and exchange mechanisms entirely around full GPU residency. The CPU serves only as a fallback path when certain features are not supported on GPUs. This architecture can reduce the design complexity compared to a retrofitted or heterogeneous design, thereby improving portability and performance.
@@
Drop-In Acceleration. Modern software stacks are gradually moving toward composable architectures (Pedreira et al., 2023), where components such as query optimizers, storage format, and query execution engine are decoupled. Sirius embraces this modular design to significantly reduce the development cost and the barrier of adoption.
@@ 3.2.3. Buffer Manager
Data Processing. This region is in device memory which stores intermediate results during query execution (e.g., hash tables, intermediate results, etc.). Sirius uses the RMM (rmm, 2025) (RAPIDS Memory Manager) pool allocator to manage this region efficiently.
@@ 3.4. Future Extensions
Out-of-Core Execution. Currently, Sirius operates under the assumption that the input data fits in the caching region and the intermediate results fit in the processing region. To support larger-than-memory workloads, we plan to extend the buffer manager to support spilling to pinned memory and disk. We will also enhance the execution engine to support batch execution by partitioning input data into batches and pipelining them into the GPU (Yuan et al., 2025; Chrysogelos et al., 2019).
Full Distributed and Multi-GPU Support. The current distributed mode offers limited SQL coverage compared to its single-node counterpart. For instance, it does not support functions such as avg. Going forward, we plan to expand the SQL coverage for the distributed execution and introduce other key features such as fault tolerance. Additionally, we will extend Sirius to support multiple GPUs per node (Yogatama et al., 2024), enabling it to handle larger workload.
Expanding SQL Coverage. Sirius already supports most of the basic SQL operators and data types. We plan to broaden the coverage by adding support for more complex data types, such as LIST, and nested types. Additionally, we aim to implement advanced SQL operators, including ASOF joins, vector search, and UDFs.
Expanding Host Database Coverage. Sirius currently serves as a drop-in accelerator for both DuckDB and Apache Doris. Going forward, we plan to support other substrait-compatible systems, such as DataFusion, Ibis, and Gluten (glu, 2025).
Performance Optimizations. Sirius is still in its early stages, and we see significant opportunities for performance improvement. Future optimizations include predicate transfer (Yang et al., 2023; Yang and Yu, 2025), efficient distributed shuffling, operator-specific enhancements (Wu et al., 2025), and lightweight compression techniques (Shanbhag et al., 2022; Afroozeh et al., 2024) to mitigate GPU memory capacity limitations—particularly during batch/pipeline execution (Boeschen et al., 2024). Additionally, we aim to optimize I/O paths for reading data from disk (Torp et al., 2025) and across the network using GPUDirect (gpu, 2025).
@@ 4.1. Experiment Setup
Benchmark. We use the TPC-H benchmark in our experiment, which has been widely used both in academia and industry.
@@
DuckDB: DuckDB (Raasveldt and Mühleisen, 2019) is a single node open-source embedded CPU-based analytical database.
ClickHouse: ClickHouse (Schulze et al., 2024) is a distributed open-source CPU-based online analytical database.
Apache Doris: Doris (dor, 2025) is an open source high-performance and real-time distributed CPU-based data warehouse.
@@
4.2. Single-Node Performance
In this Section, we evaluate Sirius performance running on a single node as a drop-in accelerator for DuckDB. We ran 22 TPC-H queries with a scale factor of 100 and compare the performance against DuckDB and ClickHouse. To ensure cost-normality, we ran DuckDB and ClickHouse on a CPU instance (m7i.16xlarge@AWS) that has the same hourly rental cost ($3.2/h) as the GPU instance (GH200@LambdaLabs) used to run Sirius.
End-to-end Performance Comparison. Figure 4 shows the end-to-end performance of various query engines on the TPC-H benchmark. Compared to DuckDB, Sirius achieves a 7 speedup with the same hardware rental cost. Sirius leverages DuckDB’s optimized logical plans but replaces its backend with GPUs, demonstrating the significant performance advantage of GPU acceleration.
Since ClickHouse does not support correlated subqueries (Schulze et al., 2024), we rewrite queries containing subquery correlation for compatibility. Compared to ClickHouse, Sirius is 20 faster with the same hardware rental cost. Q9 does not finish in ClickHouse and Q21 is not supported. We observed that ClickHouse is not optimized for join-heavy workloads, leading to suboptimal performance in many TPC-H queries (e.g., Q2, Q5, Q10, etc).
@@ 4.2. Single-Node Performance
4.3. Distributed Performance
In this section, we ran a subset of TPC-H queries (i.e., Q1, Q3, and Q6) that are currently supported by the distributed mode of Sirius and compare the performance of Sirius against Doris and ClickHouse using a cluster of 4 A100 GPUs.
@@ 5. Conclusion
This paper argues that the GPU era for data analytics has arrived and supports this claim by examining recent trends in hardware and software development. Motivated by these trends, we introduce Sirius, an open-source GPU-native SQL engine that provides drop-in acceleration across a variety of existing data systems. Our evaluation on TPC-H benchmark shows that Sirius achieves 7 and up to 12.5 speedup when integrated with DuckDB (single-node) and Doris (distributed), respectively. With a permissive Apache-2.0 license, Sirius welcomes contributions from researchers and practitioners in the database community with the shared mission of kickstarting the GPU eras for data analytics.

### v3

@@ Abstract.
The era of GPU-powered data analytics has arrived. In this paper, we argue that recent advances in hardware (e.g., larger GPU memory, faster interconnect and IO, and declining cost) and software (e.g., composable data systems and mature libraries) have removed the key barriers that have limited the wider adoption of GPU data analytics. We present Sirius, a prototype open-source GPU-native SQL engine that offers drop-in acceleration for diverse data systems. Sirius treats GPU as the primary engine and leverages libraries like libcudf for high-performance relational operators. It provides drop-in acceleration for existing databases by leveraging the standard Substrait query representation, replacing the CPU engine without changing the user-facing interface. Sirius achieves 8.3 and 7.4 better cost efficiency on TPC-H and ClickBench, respectively, when integrated with single-node DuckDB, and delivers up to 12.5 speedup when integrated with Apache Doris distributed engine.
00footnotetext: ∗Equal contributions1. Introduction
@@
| Rental Cost | $7.344/h (AWS) | $1.5/h (Lambda Labs) |
Despite their great potential, however, GPU-based SQL engines (omnisci; kinetica; theseus; cudf; sparkrapids; theseus) have not yet seen mainstream adoption. Prior efforts, both academic and commercial, have been constrained by the following four key challenges and thereby remain niche solutions.
@@
In this paper, we argue that the hardware and software foundations are finally in place today to overcome these challenges and enable scalable GPU data analytics. On the hardware side, GPU memory capacity doubles almost every generation, starting from Volta (32 GB), to Ampere (80 GB), Hopper (192 GB), and most recently Blackwell (288 GB). Better interconnects such as PCIe Gen6, NVLink-C2C (nvlinkc2c), and GPU Direct (gpudirect) significantly reduce the data movement overhead between GPU and other system components. This allows a GPU to process data beyond on-device memory with very fast speed, enabling TBs of analytics even on a single GPU and more with distribution. Meanwhile, GPUs are increasingly affordable and accessible, especially for older generations which are already sufficient for data analytics workloads.
On the software side, the GPU ecosystem has significantly matured. Libraries such as libcudf (cudf) provides high-performance primitives like joins and aggregations that a GPU SQL engine can build upon. The rise of composable data systems (composable)—driven by open standards like Substrait (substrait) for query representation, Apache Arrow and Parquet for columnar data formats—enable greater interoperability. A modern GPU engine can reuse existing components, such as SQL parser and query optimizer, to drastically reduce the complexity of building the system from the ground up.
To demonstrate the feasibility of GPU-native databases, we have built Sirius111https://github.com/sirius-db/sirius, a prototype open-source SQL engine that uses GPU as the primary execution device and delivers drop-in acceleration across diverse data systems. Sirius integrates seamlessly with existing database systems via the standard Substrait query representation. For example, plugging Sirius into DuckDB causes minimal change to the user-facing interface. Instead of executing the query using DuckDB’s CPU engine, the Substrait plan is routed to Sirius for GPU-native execution. Sirius builds on GPU libraries such as libcudf (cudf), RMM (rmm), and NCCL (nccl), reusing optimized implementations of core relational operators like joins, filters, aggregations, and data shuffle. Thanks to its modular design, Sirius also allows developers to easily switch the operator implementation between these GPU libraries and custom CUDA kernels. This flexibility reduces engineering overhead while enabling Sirius to achieve end-to-end GPU acceleration. Our detailed evaluation shows that Sirius delivers 8.3 and 7.4 better cost efficiency on TPC-H and ClickBench, respectively, when used as a drop-in accelerator for DuckDB (duckdb) (single-node), and up to 12.5 speedup as a drop-in accelerator for Apache Doris (doris) (distributed).
@@ 2.1. Hardware Trend
PCIe, the interconnect between GPU and CPU, has traditionally been a severe performance bottleneck due to its relatively low bandwidth. However, PCIe bandwidth has been doubling every two years with PCIe Gen6 offering 128 GB/s, which is comparable to CPU memory bandwidth. The NVLink-C2C (nvlinkc2c) technology takes this even further, enabling 900 GB/s unidirectional bandwidth between CPU and GPU. In a GH200 superchip, the Hopper GPU can access host memory at more than 400 GB/s, which is more than the maximum bandwidth between the CPU and main memory. These faster interconnects allow a GPU analytics engine to support data larger than the GPU device memory, enabling high-speed access to terabytes of main memory and beyond.
Faster Network and Storage. Besides faster memory speed, modern GPUs also have increasingly better support for network and storage, allowing a GPU engine to go beyond in-memory data processing. Technologies like GPUDirect Storage and GPUDirect RDMA allow GPUs to access data directly from storage or network with minimal CPU involvement at very high speed. Similar to interconnect technology, storage and network bandwidth have also grown rapidly in recent years—and are expected to continue improving (see Figure 1(c) and 1(d)). Moreover, the recent S3 over RDMA technique further enables high-speed object store for GPUs, which can achieve storage access speed at 200 GB/s (s3rdma). With significantly higher computational power, GPUs are better positioned than CPUs to fully utilize these high-bandwidth resources.
Declining GPU Cost. While GPUs of the latest generation remain in high demand with high cost, older generations are becoming increasingly accessible in the cloud at substantially lower cost. For instance, the GH200 instance shown in Table 1 has an on-demand price of $3.2/hour in Lambda Labs, which is cheaper than many high-end CPU instances. The on-demand H100 prices in many GPU cloud providers has seen significant price drop from $8/hour in March 2023 to around $3/hour in 2025 (h100-price-drop1; h100-price-drop2). Even in major cloud provider such as AWS, in 2025 alone, the A100, H100, and H200 on-demand price have dropped by 33%, 44%, and 25% respectively (aws-price-drop).
@@ 2.2. Software Trend
Composable Data Systems. Modern open-source data systems are embracing composability (composable) (e.g. Velox (velox)), where different system components can be independently developed and interoperate in a data system. A new database can reuse existing components such as the language frontend, query optimizer, or execution engine instead of reinventing the wheel. A crucial component in these systems is Substrait (substrait), a unified IR for physical and logical representations of query plans. The trend of composable data systems significantly lowers the barrier to build a GPU database—only the execution engine needs to be developed from the ground up tailored for GPU but other system components could be reused.
Maturing GPU Libraries. Open-source GPU libraries optimized for data processing have emerged and are rapidly reaching maturity in recent years. For example, libcudf (cudf) provides GPU-efficient implementation for relational operators such as joins, aggregations, and sorting. RMM (rmm) complements this by offering a high-performance GPU memory allocator, serving as a foundation for GPU-centric buffer managers. On the distributed side, the NCCL (nccl) provides high-speed communication between GPU devices over various interconnect, such as PCIe, NVLink, and Ethernet/RDMA.
@@ 3.1. Design Overview
Taking this into account, we design Sirius as a GPU-native engine: it treats GPUs as the primary execution engine, aiming to run the entire query plan—from scan to result—on the GPU. Sirius differs from systems that retrofit GPU acceleration onto traditional CPU-optimized engines while maintaining backward compatibility (sparkrapids; velox-cudf), or hybrid systems that split execution between CPU and GPU (mordred; maximus; lancelot; hetexchange). Instead, it builds its execution model, buffer management, and exchange mechanisms entirely around full GPU residency. The CPU serves only as a fallback path when certain features are not supported on GPUs. This architecture can reduce the design complexity compared to a retrofitted or heterogeneous design, thereby improving portability and performance.
@@
Drop-In Acceleration. Modern software stacks are gradually moving toward composable architectures (composable), where components such as query optimizers, storage format, and query execution engine are decoupled. Sirius embraces this modular design to significantly reduce the development cost and the barrier of adoption.
@@ 3.2.3. Buffer Manager
Data Processing. This region is in device memory which stores intermediate results during query execution (e.g., hash tables, intermediate results, etc.). Sirius uses the RMM (rmm) (RAPIDS Memory Manager) pool allocator to manage this region efficiently.
@@ 3.4. Future Extensions
Out-of-Core Execution. Currently, Sirius operates under the assumption that the input data fits in the caching region and the intermediate results fit in the processing region. To support larger-than-memory workloads, we plan to extend the buffer manager to support spilling to pinned memory and disk. We will also enhance the execution engine to support batch execution by partitioning input data into batches and pipelining them into the GPU (vortex; hetexchange).
Full Distributed and Multi-GPU Support. The current distributed mode offers limited SQL coverage compared to its single-node counterpart. For instance, it does not support functions such as avg. Going forward, we plan to expand the SQL coverage for the distributed execution (theseus) and introduce other key features such as fault tolerance. Additionally, we will extend Sirius to support multiple GPUs per node (lancelot), enabling it to handle larger workload.
Expanding SQL Coverage. Sirius already supports most of the basic SQL operators and data types. We plan to broaden the coverage by adding support for more complex data types, such as LIST, and nested types. Additionally, we aim to implement advanced SQL operators, including ASOF joins, vector search, and UDFs (gpu-udaf).
Expanding Host Database Coverage. Sirius currently serves as a drop-in accelerator for both DuckDB and Apache Doris. Going forward, we plan to support other substrait-compatible systems, such as DataFusion, Ibis, and Gluten (gluten).
Performance Optimizations. Sirius is still in its early stages, and we see significant opportunities for performance improvement. Future optimizations include predicate transfer (predicate-transfer; dist-pt), efficient distributed shuffling, operator-specific enhancements (gpu-join-eth), and lightweight compression techniques (gpu-compression; fastlane-gpu) to mitigate GPU memory capacity limitations—particularly during batch/pipeline execution (theseus; golap). Additionally, we aim to optimize I/O paths for reading data from disk and across the network using GPUDirect (gpudirect).
@@ 4.1. Experiment Setup
Benchmark. We use the TPC-H benchmark and ClickBench (clickbench) in our experiment, which has been widely used both in academia and industry.
@@
DuckDB: DuckDB (duckdb) is a single node open-source embedded CPU-based analytical database.
ClickHouse: ClickHouse (clickhouse) is a distributed open-source CPU-based online analytical database.
Umbra: Umbra (umbra) is a single-node RDBMS designed to process data at line-speed for both in-memory or storage resident data.
Apache Doris: Doris (doris) is an open source high-performance and real-time distributed CPU-based data warehouse.
@@
4.2. TPC-H Single-Node Performance
In this Section, we evaluate Sirius performance running on a single node as a drop-in accelerator for DuckDB. We ran 22 TPC-H queries with a scale factor of 100 and compare the performance against DuckDB and ClickHouse. To ensure cost-normality, we ran DuckDB and ClickHouse on a CPU instance (c8i.8xlarge@AWS) that has the same hourly rental cost ($1.5/h) as the GPU instance (GH200@LambdaLabs) used to run Sirius.
End-to-end Performance Comparison. Figure 4 shows the end-to-end performance of various query engines on the TPC-H benchmark. Compared to DuckDB, Sirius achieves a 8.3 speedup with the same hardware rental cost. Sirius leverages DuckDB’s optimized logical plans but replaces its backend with GPUs, demonstrating the significant performance advantage of GPU acceleration.
Since ClickHouse does not support correlated subqueries (clickhouse), we rewrite queries containing subquery correlation for compatibility. Compared to ClickHouse, Sirius is 38 faster with the same hardware rental cost. Q9 does not finish in ClickHouse and Q21 is not supported. We observed that ClickHouse is not optimized for join-heavy workloads, leading to suboptimal performance in many TPC-H queries (e.g., Q2, Q5, Q10, etc).
@@ 4.2. TPC-H Single-Node Performance
4.3. TPC-H Distributed Performance
In this section, we ran a subset of TPC-H queries with a scale factor of 100 (i.e., Q1, Q3, and Q6) that are currently supported by the distributed mode of Sirius and compare the performance of Sirius against Doris and ClickHouse using a cluster of 4 A100 GPUs.
@@ 4.3. TPC-H Distributed Performance
4.4. ClickBench Performance
In this section, we evaluate Sirius as a drop-in accelerator for DuckDB and compare its performance against the top two systems on the ClickBench leaderboard (Umbra and DuckDB222as of December 5, 2025; results are subject to change). We ran Sirius on a GH200 instance from Lambda Labs ($1.5/hour). The competing systems ran on CPU instances—AWS c8g.metal-48xl ($7.6/hour) and AWS c7a.metal-48xl ($9.8/hour). Because the three systems use differently priced hardware, we report the normalized total cost of ownership (TCO) relative to Sirius.
Figure 6 shows the normalized TCO of Umbra and DuckDB relative to Sirius (fixed at 1). Sirius achieves the lowest TCO on all queries, benefiting from highly efficient GPU execution. In particular, queries q4, q5, and q18 exhibit substantial performance gains on common operators such as filtering, projection, and aggregation. Some queries, however, highlight optimization opportunities. For example, q23 is bottlenecked by the contains operation on string columns; q24 and q26 by top-N operators; and q27 by aggregations over very large inputs. We plan to improve these operators in future releases of Sirius.
Overall, Sirius delivers at least a 7.4× improvement in TCO when running ClickBench as a drop-in accelerator for DuckDB. It achieves the lowest relative runtime on the ClickBench leaderboard—11% faster than Umbra, the top performer—while running on significantly cheaper hardware ($1.5/hour vs. $9.8/hour), setting a new performance record for the benchmark.
@@ 5. Conclusion
This paper argues that the GPU era for data analytics has arrived and supports this claim by examining recent trends in hardware and software development. Motivated by these trends, we introduce Sirius, an open-source GPU-native SQL engine that provides drop-in acceleration across a variety of existing data systems. Our evaluation on TPC-H and ClickBench shows that Sirius delivers 7.4–8.3 better TCO than state-of-the-art systems in a single-node setting, and up to 12.5 speedup in distributed setting. With a permissive Apache-2.0 license, Sirius welcomes contributions from researchers and practitioners in the database community with the shared mission of kickstarting the GPU eras for data analytics.
6. Acknowledgment
This work was supported (in part) by the Lambda Cloud Credits and NVIDIA Data Science Cloud Credits. We thank Ming Liu for providing hardware access for our distributed experiments. We also thank Felipe Aramburu, Amin Aramoon, Rodrigo Aramburu, Matthijs Brobbel, Johan Peltenburg, and Dhruv Vats for their tremendous contributions to Sirius.